# Data driven discovery of human mobility models

### v1

@@ 0.2 Discovering symbolic models of mobility flows
As shown in Fig. 2 and Extended Data Fig. 2, SR models achieve superior or equal accuracy to existing models of the same complexity, with existing models consistently falling on the upper-right side of the Pareto frontiers formed by the SR models (Fig. 2 a-c). Notably, at a complexity of 5, SR autonomously discovered the gravity model with exponential decay in six of the seven datasets. In the remaining case, the accuracy of the exponential decay gravity model was almost identical to that of the model on the Pareto frontier (only 1.6% higher RMSE; Supplementary Table S3). This strong convergence across diverse datasets underscores the effectiveness and parsimony of the exponential decay gravity model for representing human mobility. Consistent with previous empirical analysis[38], the distance decay parameters of the SR-derived gravity models show a decreasing trend with the area of the spatial units, from 0.21 for Guangdong sub-districts to 0.04 for US counties (Supplementary Table S1). Theoretically, Wilson’s maximum entropy derivation of the exponential decay ($1/e^{c_{ij}}$, where $c_{ij}$ is the travel cost from location $i$ to $j$) assumes travel cost is proportional to distance (see Methods). Our empirical findings provide strong support for this theoretical interpretation: by exploring a broad space of functional forms, SR identifies the exponential decay gravity model as a near-optimal solution balancing predictive power and model simplicity. The proximity of this model to the Pareto frontier’s inflection point further suggests that the exponential decay gravity model effectively captures the distribution of flow data.
@@
Beyond the exponential decay, SR discovered a novel distance decay function, $f(d_{ij})=e^{d_{ij}^{\alpha}}$, at the complexity of 7 in both the England and US commuting datasets, with $\alpha=0.50$ and $0.42$, respectively (Fig. 2d). This new form substantially improves predictive accuracy, reducing RMSE by 4.64% in the England data and 9.23% in the US data. Within Wilson’s maximum entropy framework mentioned above, this decay form corresponds to a power-law relationship between travel cost and distance ($c_{ij}\propto d_{ij}^{\alpha}$, see Methods). While a linear relationship is commonly assumed in transportation research, the relationship between travel cost (including monetary and time cost) and distance can be complicated by mixed transportation modes, leading to different function forms[39]. Our finding provides a new hypothesis for empirical validation in the future. Regardless, the improved predictive performance of this interpretable decay function highlights its potential for future applications in human mobility modeling.
@@
Regarding the other existing models, the gravity model with power-law decay is second best overall, with an average 11.2% higher RMSE than the exponential decay (Extended Data Table 1). Models based on intervening opportunity do not perform well on most datasets. On average, the RMSE of the radiation model is 92.0% higher than the SR model with the same complexity, and the error increase for Schneider’s intervening opportunity model and opportunity priority selection model are 39.8% and 49.1%, respectively. Besides, the intervening opportunity appears in the Pareto optimal SR models on only one of the seven datasets. Hence, we found no evidence that the intervening opportunity better describes the decay effect than geographic distance.
Fig. 3: Spatial heterogeneity of the mobility model across US. (a) The distance distribution of commuting flows, grouped by geographic regions. The predicted flows are from the complexity 5 SR model on each subset grouped by the origin and destination region. For inter-region flows, each subplot shows outflows from one region, and the line color corresponds to the destination region. (b) SR models at complexity 5 for commuting flows. Each row/column represents the region containing the residential/work place. The notations in formulas are the same as in Fig. 1. The accuracy of these models is compared with one of the power-law and exponential gravity models which produces lower error on the subset of flow data. A red/blue cell indicates the accuracy of the SR model is better/worse (measured with RMSE); in such cases, the gravity model is given in parentheses. A grey cell means the SR model is identical to the gravity model.
@@ 0.3 Geographical heterogeneity of mobility models
For all four subsets of intra-region flows, the exponential decay gravity model remains optimal as suggested by SR (the diagonal expressions in Fig. 3b). This provides the basis for the model’s overall optimality as intra-region flows dominate commuting flows, containing 97.5% of flow volumes. Yet the exponential decay gravity model appears to be suboptimal for inter-region commuting. Over the 12 subsets of inter-region flows, the power-law decay gravity model consistently outperforms exponential decay. Moreover, even better models are discovered by SR in 7 of the 12 subsets.
[…]
@@ 0.4 Stability of Symbolic Regression under noise
All observational data contain noise, and if the noise is high enough, it is likely that we will not be able to identify valid information or models[43]. To evaluate the effectiveness of SR in the face of noisy mobility data, we conducted experiments on synthetic data (since the noise distribution in the real data is unknown). Specifically, we align the settings with our analysis on US data. Using the real spatial boundary, workplace population, and total outflow, we generate origin-destination flows based on three mobility models: the exponential decay gravity model, the power-law decay gravity model, and radiation model. We generated five samples of flow data at each of 9 noise levels (additive Gaussian noise with the standard error $\sigma$ increasing from 0 to 2.0 with a step of 0.25), where the middle level ($\sigma=1.0$) is comparable to the amount of random noise in real data (see Methods). The SR program was independently applied to each sample.
@@
Figure 4 reports the number of successful instances where SR discovers the correct form of the generation model. When no random noise is applied, all three models can be successfully distilled. Furthermore, under a moderate amount of noise (below $\sigma=1.0$), SR still discovers the two gravity models with 100% success rate, while the discovery of the radiation model becomes unstable as the noise level increases (Fig. 4). The latter, however, does not imply that SR produces sub-optimal models. In most cases, the SR models with complexity 11 actually have better accuracy than the radiation model, which indicates the model that best describes the data has changed because of the noise introduced. Such “shifted” formulas discovered by SR are inconsistent across samples. Some of them resembles the radiation models, such as $w_{j}/(w_{i}+s_{w}+aw_{j})^{\beta}$, $w_{j}^{\alpha}/(w_{i}+s_{w}+a)^{\beta}$, $w_{j}/(w_{i}+s_{w})(w_{i}+as_{w})$, where $a,\alpha,\beta$ are constants. These results suggest ways to extend the radiation model by introducing parameters to better model noisy data.
Fig. 4: The success rate of SR to reproduce generation models on simulated data. The additive Gaussian noise on the logarithm of flows is applied, with the minimum noise level to exceed real data (measured by model CPC) marked. Symbolic regression remains robust to random noise. Only at relatively high levels of noise do three generation models fail to be successfully discovered by symbolic regression.
Discussion
Automatic distilling symbolic models from social data can be more challenging than from physical data, as human behavior is difficult to measure with explicit expressions. In this work, we show that SR is capable of discovering accurate and compact models that capture the regularities in human mobility, and the derived Pareto optimal models reveal feasible ways to extend the basic gravity formula, some of which are previously unknown. More generally, the accumulation of observational data and the progress in symbolic regression provide a promising way to build empirical models in the social sciences. Rather than comparing the accuracy of existing candidate models, SR allows us to answer the questions: Does an existing model have the best form among all possible expressions, given a set of operators? If not, does SR suggest new candidate models of theoretical interest? In the future, it is intriguing to apply SR to more location features other than population for mobility model discovery. We also look forward to applications of SR beyond mobility data, which may provide new perspectives on the quantitative modeling of social phenomena.
@@ Discussion
We gain insight into human mobility from not only the SR-derived explicit mathematical models, but also how key factors are progressively incorporated as the allowed model complexity increases. As shown in Fig. 2d and Extended Data Fig. 2c, SR first captures the distance friction effect——near destinations attract more people than distant ones——at complexity 3, suggesting it is the principal effect governing human mobility. Another key factor revealed by SR is the destination attractiveness (represented by population), which is mostly incorporated into the models at complexity 5. Places with larger populations tend to attract more mobility flows, as they provide a greater number and diversity of goods, services, and job opportunities. Most SR models at complexity 7 and 9 share the form of $g(w_{j},r_{j})/f(d_{ij})$, which still combine the destination attractiveness and the distance friction effect, yet with more complex function forms for one or both effects. As complexity further increases, the SR models become inconsistent across datasets and difficult to explain. These models are not likely to be general models for mobility flows, but capturing the characteristics of a single dataset.
@@
Our analysis also reveals the limitations of Genetic Programming-based SR. Although the performance of SR on simple formulas (complexity$<$9) is generally satisfactory, the derived formulas for longer expressions are inconsistent across repeated runs, indicating that sub-optimal solutions are found. This may be due to the exponentially growing search space. Existing literature also points out that SR is only robustly applicable to datasets with fewer than 10 variables[25]. Incorporating neural networks and symbolic regression might be a feasible approach to discover more complicated models. For example, SR could be applied to separable modules of a neural network, and the distilled formulas are then combined into a complete expression[44, 45, 29]. Future studies could investigate the effectiveness of such neuro-symbolic approaches for knowledge discovery in the social sciences.
Methods
0.5 Empirical data.
The Guangdong dataset is provided by China Mobile Communications Corporation, the largest telecommunications operator in China. The raw dataset comprises individual movement trajectories of 5 million users in November 2020, with a spatial resolution of 500 m $\times$ 500 m. For the mobility flow dataset, individual movements are aggregated at the sub-district and county levels based on the grids’ center locations. For the commuting flow dataset, an individual’s home is defined as the location where they stayed the longest between 12:00 AM and 7:00 AM, while the workplace is defined as the location where they stayed the longest between 9:00 AM and 6:00 PM. Commuting flows are identified based on movements from the home location to the workplace in the morning (before 12:00 PM). Subsequently, commuting movements are aggregated similarly to the mobility flow dataset. The residential / workplace populations are derived based on the number of residents and workers within each spatial unit.
The Beijing-Tianjin-Hebei mobility data is provided by China Unicom, which contains aggregated inter-county mobility from November 4 to 10, 2019. The population data in 2019 are collected from official statistical yearbooks.
The commuting datasets in England and the Contiguous US record the location of usual residence and place of work of employed residents, aggregated to MLAD and county levels, respectively. The sources of residential and workplace population data are the same as the flow data (2011 Census for England, and 2011-2015 ACS 5-year estimate for US).
For all datasets above, we only consider flows between spatial units, so flows with the same origin and destination are removed. Moreover, the intervening opportunity is based on the population data at finer scale [Middle layer Super Output Area (MSOA) in England, sub-district in Guangdong]. The number and size of spatial units, as well as the number of observations (origin-destination pairs) of each dataset, are given in Supplementary Table S1. The mobility flows are visualized in Extended Data Figs. 3-7. For calibration of existing models, workplace populations are used on commuting datasets for better accuracy than residential populations, while only residential populations are available for mobility datasets.
0.6 Simulated data.
We introduce random noise on model-generated flows to represent factors of mobility which are not captured by the symbolic models. We apply additive Gaussian noise on the logarithm of flows, which ensures the flows are non-negative. Assume $\bar{T}_{ij}$ is the model-generated flow from place $i$ to $j$, the synthetic flow $T_{ij}$ is sampled as
$T_{ij}=\bar{T}_{ij}\exp(\epsilon),\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (1)
where $\mathcal{N}(0,\sigma^{2})$ is the Gaussian distribution with mean zero and variance $\sigma^{2}$. We do not consider a simple additive Gaussian noise for potential heteroskedasticity, as the variation of large and small flows are not likely to be the same. To align with real data, we rounded $T_{ij}$ to the closest integer. Flows less than 3 are removed for symbolic regression.
To compare with the potential noise level in the real data, we use Common Part of Commuters (CPC)[46] to measure the discrepancy between the model-predicted flows and real flows. As a common metric for flow prediction, CPC is defined as:
$\text{CPC}=\frac{2\sum_{i,j}\min\{F_{ij},\hat{F}_{ij}\}}{\sum_{i,j}F_{ij}+\sum_{i,j}\hat{F}_{ij}}$ (2)
where $F_{ij},\hat{F}_{ij}$ are the real and predicted flow volume from place $i$ to $j$, respectively. The range of CPC is $[0,1]$, and lower value indicates a worse match between real and predicted flows, hence higher level of random fluctuation. Using CPC of the exponential decay gravity model on US data (0.728) as a threshold, we identify the minimum noise level in simulated data where the CPC of the generation model falls under 0.728, which is $\sigma=1.0$ for all three considered generation models (Fig. 4).
0.7 Symbolic Regression.
Given a set of functions $\mathcal{P}$, a dataset $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$. Assume :×$\mathcal{L}:\mathbb{R}\times\mathbb{R}\to[0,\infty)$ is a distance function, where common choices are squared error or absolute error; $C$ is a measure of function complexity. Symbolic regression (SR) may be formulated as a minimization problem to find
$f^{*}\in\operatorname{\arg\min}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(y_{i},f(\mathbf{x}_{i}))+\lambda C(f)$ (3)
where $\mathcal{F}$ is the set of functions which can be formed by composition of the functions in $\mathcal{P}$; $\lambda$ is a penalty factor for complexity. In practice, SR is often solved as a multi-objective optimization problem which jointly minimizes the fitting error and the model complexity[20]. SR is proven to be NP-hard[47], which indicates an exact and efficient solution might not exist. Though methods based on physical principles[48, 49], neural networks[50, 51, 52], and Bayesian computation[53, 54] are emerging recently, genetic programming[55, 56, 32] remains the mainstream approach to SR.
We utilize the Julia package SymbolicRegression.jl[20] as it is specially designed for symbolic regression on scientific data. Several enhancements to the general framework of genetic programming are implemented, including the integration of simulated annealing, age-regularization to avoid premature convergence, expression simplification and constant optimization, a modified loss to balance expression optimization at all complexity, and migration of expressions between several parallel populations. The output contains all Pareto optimal expressions regarding accuracy and complexity. We refer to [20] for technical details. Nevertheless, optimization of the allocation weight function is not readily supported by the package. We develop our SR program by modifying the evaluation function in SymbolicRegression.jl to calculate flow volumes from allocation weights. As constant optimization only takes place with probability in the genetic search process, we also add a constant optimization step before the output of final expressions.
Due to the randomness of the genetic algorithm, we repeat the SR procedure multiple times for each dataset. Typically, we first use all features as explanatory variables, then restrict the set of variables to those shown to be important. Fewer variables lead to a smaller search space, thus may yield better solutions. The number of iterations is set as 100 throughout this study. Batching is necessary to accelerate the program for large flow datasets. By default, the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is used to optimize constants, and the Nelder-Mead algorithm is used if the former fails. The Pareto optimal expressions within the combined output of repeated runs are selected as final models. The detail of SR settings on each dataset is given in Supplementary Table S2.
0.8 Maximum entropy approach to the gravity model.

### v2

@@ 3.1 Discovering symbolic models of mobility flows
As shown in Fig. 3 and Supplementary Fig. S1, SR models achieve superior or equal accuracy to existing models of the same complexity, with existing models consistently falling on the upper-right side of the Pareto frontiers formed by the SR models (Fig. 3 a-c). Notably, at a complexity of 5, SR autonomously discovered the gravity model with exponential decay in six of the seven datasets. In the remaining case, the accuracy of the exponential decay gravity model was almost identical to that of the model on the Pareto frontier (only 1.6% higher RMSE; see Supplementary Table S2). This strong convergence across diverse datasets underscores the effectiveness and parsimony of the exponential decay gravity model for representing human mobility. Consistent with previous empirical analysis [22], the distance decay parameters of the SR-derived gravity models show a decreasing trend with the area of the spatial units, from 0.21 for Guangdong sub-districts to 0.04 for US counties (Table 1). Theoretically, Wilson’s maximum entropy derivation of the exponential decay ($1/e^{c_{ij}}$, where $c_{ij}$ is the travel cost from location $i$ to $j$) assumes travel cost is proportional to distance (see Appendix). Our empirical findings provide strong support for this theoretical interpretation: by exploring a broad space of functional forms, SR identifies the exponential decay gravity model as a near-optimal solution balancing predictive power and model simplicity. The proximity of this model to the Pareto frontier’s inflection point further suggests that the exponential decay gravity model effectively captures the distribution of flow data.
Figure 3: SR results on mobility flow data. (a-c) Pareto frontiers of SR models on Guangdong, England, and US datasets. As flow magnitudes vary across datasets, we normalize the RMSE with that of the simplest gravity model ($w_{j}/d_{ij}$). The accuracy and complexity of six existing models are marked with crosses (note that some existing models are not shown as their errors exceed the range of the y-axis). Expressions with complexity 1 are excluded as one variable or constant is not sufficient to model the mobility flow. The expressions with complexity 15 on the US dataset is not Pareto optimal. Across all datasets and complexity levels, the SR models consistently outperform or match the existing models in accuracy. (d) Expressions of the Pareto optimal SR models. The notations follow Fig. 1, except that $d$ is short for $d_{ij}$. To align with existing models, some expressions are not given in the form with the lowest complexity. The distilled models are classified based on the captured effects governing human mobility. (e) Expressions of the considered existing models.
@@
Beyond the exponential decay, SR discovered a novel distance decay function, $f(d_{ij})=e^{d_{ij}^{\alpha}}$, at the complexity of 7 in both the England and US commuting datasets, with $\alpha=0.50$ and $0.42$, respectively (Fig. 3d). This new form substantially improves predictive accuracy, reducing RMSE by 4.6% in the England data and 9.3% in the US data. Within Wilson’s maximum entropy framework mentioned above, this decay form corresponds to a power-law relationship between travel cost and distance ($c_{ij}\propto d_{ij}^{\alpha}$, see Appendix). While a linear relationship is commonly assumed in transportation research, the relationship between travel cost (including monetary and time cost) and distance can be complicated by mixed transportation modes, leading to different function forms [59]. Our finding provides a new hypothesis for empirical validation in the future. Regardless, the improved predictive performance of this interpretable decay function highlights its potential for future applications in human mobility modeling.
@@
Regarding the other existing models, the gravity model with power-law decay is second best overall, with an average 11.2% higher RMSE than the exponential decay (Table 2). Models based on intervening opportunity do not perform well on most datasets. On average, the RMSE of the radiation model is 92.0% higher than the SR model with the same complexity, and the error increase for Schneider’s intervening opportunity model and opportunity priority selection model are 39.8% and 49.1%, respectively. Besides, the intervening opportunity appears in the Pareto optimal SR models on only one of the seven datasets. Hence, we found no evidence that the intervening opportunity better describes the decay effect than geographic distance.
Table 2: The accuracy of existing models on flow data, measured by root mean squared error (RMSE). The lowest error on each dataset is put in bold. See Fig. 3e for model definitions. BTH is short for Beijing-Tianjin-Hebei urban agglomeration, China.
Type Commuting Mobility
Region Guangdong England US Guangdong BTH
Granularity sub-district county MLAD county sub-district county county
GMZipf 15.02 146.25 728.58 2105.9 1588.5 12870 81813
GMPow 12.53 123.78 442.23 1337.7 1173.6 11734 76959
GMExp 11.51 111.91 360.34 1309.9 1112.8 10723 64746
RM 22.72 105.62 863.67 2222.4 2023.1 14628 119256
IO 14.20 108.49 587.86 1733.4 1289.8 12260 74732
OPS 12.99 131.69 679.65 2370.0 1332.9 13582 85542
@@ 3.2 Geographical heterogeneity of mobility models
For all four subsets of intra-region flows, the exponential decay gravity model remains optimal as suggested by SR (the diagonal expressions in Fig. 4b). This provides the basis for the model’s overall optimality as intra-region flows dominate commuting flows, containing 97.5% of flow volumes. Yet the exponential decay gravity model appears to be suboptimal for inter-region commuting. Over the 12 subsets of inter-region flows, the power-law decay gravity model consistently outperforms exponential decay. Moreover, even better models are discovered by SR in 8 of the 12 subsets.
Figure 4: Spatial heterogeneity of the mobility model across US. (a) The distance distribution of commuting flows, grouped by geographic regions. The predicted flows are from the complexity 5 SR model on each subset grouped by the origin and destination region. For inter-region flows, each subplot shows outflows from a region, and the line color corresponds to the destination region. (b) SR models at complexity 5 for commuting flows (we report the complexity 3 model for flows from West to Midwest as the complexity 5 model is not Pareto optimal). Each row/column represents the region containing the residential/work place. The notations in formulas are the same as in Fig. 1, except that $d$ is short for $d_{ij}$; $s_{r},s_{w}$ are short for $s^{(r)}_{ij},s^{(w)}_{ij}$, respectively. The accuracy of these models is compared with one of the power-law and exponential gravity models which produces lower error on the subset of flow data. A red/blue cell indicates the accuracy of the SR model is better/worse (measured with RMSE); in such cases, the gravity model is given in parentheses. A grey cell means the SR model is identical to the gravity model.
This disparity in optimal model form may indicate difference in the mechanisms governing inter-region and intra-region commuting. Actually, the distance distribution of intra-region flows follow a typical distance decay pattern, while uncommon bimodality is observed for inter-region subsets, with a long-range peak at around 1000 km (Fig. 4a). Through examination of the flow data, we find the destinations of these flows are mostly large cities, with the top 5 destinations as Chicago, New York, Los Angeles, Houston, and Dallas. We speculate that, with more job opportunities, the attractiveness of these large cities may be much stronger than the distance friction effect in long-distance commuting, leading to weak or even disappeared distance decay. This difference is also reflected in some SR models for inter-region commuting, where flow volumes do not decrease with distance as commonly assumed (e.g. models for flows between Midwest and West in Fig. 4b).
@@ 3.3 Stability of Symbolic Regression under Noise
All observational data contain noise, and if the noise is high enough, it is likely that we will not be able to identify valid information or models [12]. To evaluate the effectiveness of SR on noisy mobility data, we conducted experiments on synthetic data (since the generation model and noise structure in the real data are unknown). We align the settings with our analysis on US data. Using the real spatial boundary, workplace population, and total outflow, we generate the expected values of origin-destination flows $\mu_{ij}$ based on allocation weights determined by three mobility models: the exponential decay gravity model, the power-law decay gravity model, and radiation model (Eq.4, Fig.3e). We then apply a negative binomial model to represent random fluctuations, or factors not captured by the symbolic models [61]. Although the Poisson distribution is more commonly used in flow modeling [43], its variance is determined by the mean, while the negative binomial distribution with two parameters allows us to adjust the amount of random noise. Specifically, the probability mass function is given by
$P(T_{ij}=k)=\frac{\Gamma(k+\rho_{ij})}{k!\Gamma(\rho_{ij})}\left(\frac{\mu_{ij}}{\rho_{ij}+\mu_{ij}}\right)^{k}\left(\frac{\rho_{ij}}{\rho_{ij}+\mu_{ij}}\right)^{\rho_{ij}},$ (8)
where $\mu_{ij},\rho_{ij}$ are parameters. It follows that
$\mathrm{E}(T_{ij})=\mu_{ij},\mathrm{Var}(T_{ij})=\mu_{ij}+\frac{\mu_{ij}^{2}}{\rho_{ij}}.$ (9)
We denote $\alpha_{ij}=\frac{1}{\rho_{ij}}$, and use the same value $\alpha_{ij}=\alpha$ for all origin-destination pairs. Hence, $\mathrm{Var}(T_{ij})=\mu_{ij}+\alpha\mu_{ij}^{2}$, so higher $\alpha$ indicates larger random noise. We consider a series of $\alpha$ values from 0 to 2.0 with a step of 0.25 (we use $\alpha=10^{-8}$ as the minimal level as $\alpha=0$ is not allowed). For each $\alpha$ value, we sample each flow $T_{ij}$ independently from the negative binomial model. Finally, we rounded $T_{ij}$ to the closest integer. We generated five samples of flow data for each combination of 9 $\alpha$ values and three mobility models. The SR program was independently applied to each sample (see Table S1 for details), where we dropped flows less than 3 due to computational burden from a large number of such small flows.
To illustrate the amount of noise in the synthetic data, we apply CPC to measure the discrepancy between flows without noise $\{\mu_{ij}\}$ and flows with noise $\{T_{ij}\}$. We use the CPC value of the exponential decay gravity model (the best-fit baseline model) on the real US commuting flows, 0.728, as a standard, reflecting the potential amount of random fluctuation in the real data. For all three considered models, the synthetic data reach a comparable CPC at $\alpha=0.5$ (Supplementary Table S8).
Figure 5: The success rate of SR to identify generation models on simulated data. The negative binomial distribution is applied to model the random fluctuations in mobility flows. The noise level comparable to that in real data (measured by model CPC) is marked with dashed lines.
@@
Figure 5 reports the number of successful instances where SR discovers the correct form of the generation model. When little random noise is applied, all three models can be successfully distilled. Furthermore, under a moderate amount of noise (more than $\sigma=0.5$), SR still discovers the two gravity models in most cases, while the discovery of the radiation model becomes unstable as the noise level increases. The latter, however, does not imply that SR produces sub-optimal models. In all but one among the 25 cases where the SR model with complexity 11 is inconsistent with the radiation model, the SR models actually have lower RMSE. This indicates the model that best describes the data has changed because of the noise introduced. Although such “shifted” formulas discovered by SR are inconsistent across samples, some of them resembles the radiation model, such as $w_{j}/((w_{i}+s_{w})^{\beta}+w_{j}^{\gamma})$, $w_{j}/((w_{i}+s_{w})^{\beta}(w_{i}+w_{j}))$, $w_{j}/(w_{i}+s_{w}+w_{j}/w_{i})^{\beta}$, where $\beta,\gamma$ are constants. These formulas suggest potential ways to extend the radiation model to better model noisy data.
4 Discussion and Conclusions
@@
Genetic programming-based SR not only enables us to examine the optimality of existing models, but also sheds light on how key factors are progressively incorporated as the allowed model complexity increases. As shown in Fig. 3d and Supplementary Fig. S1c, SR first captures the distance friction effect——near destinations attract more people than distant ones——at complexity 3, suggesting it is the primary effect governing human mobility. Another key factor revealed by SR is the destination attractiveness (represented by population), which is mostly incorporated into the models at complexity 5. Places with larger populations tend to attract more mobility flows, as they provide a greater number and diversity of goods, services, and job opportunities. Most SR models at complexity 7 and 9 share the form of $g(w_{j},r_{j})/f(d_{ij})$, which still combine the destination attractiveness and the distance friction effect, yet with more complex function forms for one or both effects. As complexity further increases, the SR models become inconsistent across datasets and difficult to explain. These models are not likely to be general models for mobility flows, but capturing the characteristics of a single dataset.
Recently, [6] applied Bayesian Machine Scientist (BMS) [14], a Bayesian-based approach to Symbolic Regression, to analyze human mobility data in the US. Their study demonstrated that models derived from BMS are at least comparable to complex machine learning models in predicting mobility flows, with function forms resembling the gravity model. For comparison, we applied the BMS approach to the US and England commuting data, and report the resulting formulas in Supplementary Table S9. BMS tends to generate models of complexity over 20, while our genetic programming-based approach identifies optimized models across a range of complexity from 3 to 15, including those comparable to existing mobility models. In this sense, the two approaches can complement each other in exploring the accuracy-complexity Pareto frontier of symbolic models. Furthermore, by examining the complexity level at which each variable is included, our approach can identify the primary and secondary effects in the process under investigation, which is not feasible with BMS.
@@
Our analysis also reveals the limitation of genetic programming-based SR. Although the performance of SR on simple formulas (complexity<<9) is generally satisfactory, the derived formulas for longer expressions are inconsistent across repeated runs, indicating that sub-optimal solutions are found. Existing literature also points out that SR is only robustly applicable to datasets with fewer than 10 variables [54], due to the exponentially growing search space. Incorporating neural networks and symbolic regression might be a feasible approach to discover more complicated models. For example, SR could be applied to separable modules of a neural network, and the distilled formulas are then combined into a complete expression [10, 42, 27]. Future studies could investigate the effectiveness of such neuro-symbolic approaches for model building in human mobility.
In conclusion, this work demonstrates that SR is capable of discovering accurate and compact models that capture the regularities in human mobility, and the derived Pareto optimal models reveal feasible ways to extend the basic gravity model, some of which are previously unknown. In the future, it is intriguing to apply SR to more location features other than population for mobility model discovery. More generally, the accumulation of observational data and the progress in symbolic regression provide a promising way to build empirical models in the social sciences. Rather than comparing the accuracy of existing candidate models, SR allows us to answer the questions: Does an existing model have the best form among all possible expressions, given a set of operators? If not, does SR suggest new candidate models of theoretical interest? We look forward to applications of SR beyond mobility data, which may provide new perspectives on the quantitative modeling of social phenomena.
Acknowledgments
We thank Tianyou Cheng and Chengling Tang for assistance in processing the cellphone datasets; Andre Python, Ryan Q. Wang, Ce Hou, and Xuechen Wang for helpful discussion and comments.
Funding
We acknowledge the support of the National Natural Science Foundation of China under Grant Nos. 42422110 and 42430106. L.D. was supported by the Fundamental Research Funds for the Central Universities, Peking University.
Declaration of competing interest
The authors declare no competing interests.
Author contributions
L.D., Y.L., and H.G. conceived the project and designed the experiments; H.G., W.Z., and Y.H. collected the data; H.G., W.Z., and J.Y. performed the experiments; W.Z., and H.G., visualized the results; H.G., L.D., W.Z, and Y.L. wrote the manuscript.
Data and code availability
All codes and processed flow datasets from the US, the UK, and Beijing-Tianjin-Hebei, China are publicly available on GitHub at https://github.com/urbansci/FlowSR. The Guangdong cellphone dataset can be requested from the corresponding author.
Appendix A Maximum entropy approach to the gravity model